OpenStudy (shelby1290):

Alison is at a baseball game with herbrother, John. She has $18 to spend on hamburgers and pop. Hamburgers cost $3 each and pop cost $2 each. The equation 2x + 3y=18 can be used to model this problem. A) what does the x in the equation represent? b) what does the y in the equation represent? C) graph the equation.

OpenStudy (anonymous):

The real question is.... Was Alison ever at the baseball game? Or was it her twin sister! Ha!

OpenStudy (a1234):

X represents the pops and y represents the hamburgers. To graph the equation, put it in slope-intercept form: 2x + 3y=18 3y = -2x + 18 y = -2/3x + 6

OpenStudy (shelby1290):

Loooool

OpenStudy (anonymous):

You gotta think outside the box sometimes..

OpenStudy (shelby1290):

@a1234 how do i do that?

OpenStudy (a1234):

Do what, the graphing?

OpenStudy (shelby1290):

yes

OpenStudy (anonymous):

No, investigate weather or not she was at the game. You never know

OpenStudy (a1234):

Start by locating the point (0,6), the y-intercept. Then go down two and right three. Continue to move along the graph like this. When you have enough points, draw a line through them.

OpenStudy (shelby1290):

okay

OpenStudy (shelby1290):

wait ..so, id do 1,6

OpenStudy (shelby1290):

2,6

OpenStudy (shelby1290):

and so on?

OpenStudy (shelby1290):

so all the points would be 3,4 ...4,4....and 5,4?

OpenStudy (a1234):

Your first point is (0,6), okay?

OpenStudy (shelby1290):

ok

OpenStudy (a1234):

Then go down 2 points, you will come to (0,4). But wait, DON'T plot yet! Go 3 points right from (0,4) to come to (3,4). Plot (3,4). Continue this process.

OpenStudy (shelby1290):

okay i get that

OpenStudy (a1234):

Glad you get that :).

OpenStudy (shelby1290):

where do i draw the line

OpenStudy (a1234):

Through those points.

OpenStudy (shelby1290):

like connecting the dots with an arrow

OpenStudy (a1234):

Yeah, kind of.

OpenStudy (shelby1290):

but the arrow heading towards the bottom right corner right? since its decreasing

OpenStudy (a1234):

Keep your line in first quadrant, as the numbers can only be positive. So I think you shouldn't really draw an arrow, but something like this:

OpenStudy (shelby1290):

alright thanks for the help

OpenStudy (a1234):

You're welcome.